OpenStudy (coconut):

What is the 7th term of the geometric sequence where a1 = 625 and a3 = 25?

OpenStudy (anonymous):

you've not provided enough information. What's the formula for the geometric sequence?

OpenStudy (coconut):

idk the formula :(

OpenStudy (anonymous):

a3 = sqrt of a1 a5 = sqrt of a3 a7 = sqrt of a5

OpenStudy (coconut):

my options are 0.0016 0.008 0.04 0.2

OpenStudy (anonymous):

@SomeoneYouUsedToKnow is that a geometric sequence? It's been a while for me, but I think a geometric sequence must have each successive term as a constant multiple of the previous term... So I don't think that taking square roots is an option... I might be wrong...

OpenStudy (anonymous):

It's a good guess of me because 25² = 625 thats the only hidden clue i see. In my opinion, i am correct. And yes its a geometric sequence.

OpenStudy (anonymous):

Oh yeah, my mistake.

OpenStudy (anonymous):

Happens :)

OpenStudy (anonymous):

In a geometric series, the next element is obtained by multplying the previous by a constant. So a2=c*625. Then a3=c^2*625=25. This tells you that c=1/5. Then a3=25. So a7=c^4*25=(1/5)^4*25=0.04

OpenStudy (anonymous):

For that last line, a3=25 so a4=c*25 and a5=c^2*25 and a6=c^3*25 and a7=c^4*25.

OpenStudy (anonymous):

the thing is that you said in the statement that a1 = 625 and a3 = 25. so it's wrong what you try to explain :/

OpenStudy (anonymous):

\[ r= \frac 1 5\\ a_7 = a_1 r^6= \frac 1 {25} \]

OpenStudy (anonymous):

yeah the distance between two articulars is 0,2 thats correct :)

OpenStudy (anonymous):

@SomeoneYouUsedToKnow are you referring to what I said? I don't understand...

OpenStudy (anonymous):

you just misspelled one number. you are right

OpenStudy (anonymous):

I don't think I did.